OpenStudy (anonymous):

What is the shortest distance from (0,0) to y=-4/x ?

OpenStudy (aravindg):

Hi there ! welcome to openstudy !

OpenStudy (aravindg):

About your question first convert the line into standard form

OpenStudy (anonymous):

so how would I do that?

OpenStudy (anonymous):

would someone help me please?

OpenStudy (raden):

use the formula to get distance of two points A(x1,y1) and B(x2,y2) d = sqrt{(x1-x2)^2 + (y1-y2)^2}

OpenStudy (raden):

let A=(0,0) and B(x, -4/x) apply that points to formula above

OpenStudy (raden):

then derive, just take the first derivative and make equal 0 solve for x

OpenStudy (anonymous):

why do you have to let A=(0,0) and B=(x, -4/x) ?

OpenStudy (raden):

that's ur point given, right

OpenStudy (raden):

ur question : What is the shortest distance from (0,0) to y=-4/x we need 2 point to calculate the distance of both

OpenStudy (anonymous):

okay I think I understand why. So everytime I encounter a problem like this, do I just let the point be A and the other be the y-value of another point?

OpenStudy (raden):

yeah, that's i mean :)

OpenStudy (anonymous):

okay thank you. So I just have to simply substitute them in the equation that you mentioned above and solve, right?

OpenStudy (raden):

yup, have u learned derivative before ?

OpenStudy (anonymous):

yes, I am in Calculus II.

OpenStudy (raden):

okay, i sure u can solve this

OpenStudy (anonymous):

take the derivative of the distance, right?

OpenStudy (raden):

yes

OpenStudy (anonymous):

\[d=\sqrt{x^2{+}\frac{ 16 }{ x^2{} }}\] Is that correct?

OpenStudy (raden):

yup, now derive d/dx (d')

OpenStudy (anonymous):

so \[x^4=32\]

OpenStudy (raden):

hint :derivative of sqrt(f(x)) is f ' /2sqrt(f(x))

OpenStudy (anonymous):

and x=\[2\sqrt{2}\]

OpenStudy (raden):

hmmm .. i got diferent like yours

OpenStudy (raden):

i got an equation x^4 = 16, and finally get x= +-2

OpenStudy (anonymous):

I just double checked my work and I don't think I made a mistake...

OpenStudy (raden):